### 21 Feb 2024 v3

# MoD-SLAM: Monocular Dense Mapping for Unbounded 3D Scene Reconstruction

Abstract
Neural implicit representations have recently been demonstrated in many fields including Simultaneous Localization And Mapping (SLAM). Current neural SLAM can achieve ideal results in reconstructing bounded scenes, but this relies on the input of RGB-D images. NeRF-based SLAM with RGB image input is unable to reconstruct the scale of the scene accurately, and it also suffers from scale drift due to errors accumulated during tracking. To address these issues, we present MoD-SLAM, a monocular dense mapping method that allows global pose optimization and 3D reconstruction in real-time in unbounded scenes. Optimizing scene reconstruction by monocular depth estimation and using loop closure detection to update camera pose enable detailed and precise reconstruction on large scenes. Compared to recent work, our approach is more robust, scalable and versatile. Our experiments demonstrate that MoD-SLAM has the state-of-the-art monocular mapping performance, especially in large borderless scenes.
Keywords:
3D Scene Reconstruction, Neural Radiance Field, SLAM, Monocular Depth Estimation, Unbounded Scene
1 Introduction
Dense visual Simultaneous Localization and Mapping (SLAM) is core in 3D computer vision, finding utility across various fields including autonomous driving, unmanned aerial vehicle, AR, etc. To make SLAM valuable in real-world applications, we want SLAM to be real-time, fast, accurate, and adaptable to scenes of different scales. Although traditional SLAM methods [4, 17, 47] can build maps in real-time, they can only generate sparse maps. Some works [31, 44, 47] can generate dense maps to a certain extent, but both of them are generally too slow to build maps and can not be applied to large unbounded scenes. These constraints limit the application of SLAM.
In recent years, due to the development and advancement of Neural Radiance Field (NeRF) [15, 8] and view synthesis [19, 26, 1] in 3D scene reconstruction, dense SLAM mapping has become feasible. Early NeRF-based SLAM systems such as iMAP [30] and NICE-SLAM [46] have achieved ideal results in dense mapping and tracking. However, they all have the problems of decreasing accuracy and scale drift in large scenes. Subsequent researches [42, 43, 38, 6, 25] apply the signed distance function (SDF) [22] to reconstruct the 3D scene more accurately. Nevertheless, these approaches still can not be implemented in unbounded scenes. Furthermore, numerous works rely on RGB-D input, which fails to adequately reconstruct the scale of the scene, the texture structure, and the details of geometric objects in the monocular mode. These problems limit the application and performance of neural-based SLAM systems in large scenes or scenes with numerous objects and texture details. GO-SLAM [43] inputs RGB images for 3D scene reconstruction, but the result is unsatisfactory. It lacks the prior depth maps and thus, cannot properly reconstruct the scale of the scene (Fig.1). In addition, numerous previous works lack a loop closure detection, which results in accumulated positional drift, leading to inaccuracies in tracking and mapping process.
Figure 1: Reconstruction results on ScanNet 0207 [7]. We present MoD-SLAM, a neural-based monocular dense mapping method. Our method shows a more powerful capability than existing work: GO-SLAM [43].
In this work, we propose MoD-SLAM, a monocular implicit neural-based dense mapping method. In contrast to previous works, our SLAM system no longer relies on RGB-D image inputs. We improve its accuracy by incorporating a monocular depth estimation module and a depth distillation module at the backend. These two modules allow the system to generate prior depth maps from the acquired RGB images for supervised 3D scene reconstruction. In addition, we use the Gaussian encoding that approximates the conic truncation corresponding to the pixel as input to MLPs [35] to allow the system to adapt and capture spatial information more precisely. Moreover, we present reparameterization, which enables our SLAM system to achieve ideal reconstruction results in unbounded scenes. We also include loop closure detection in the global update so that the system can combat scale drift. Extensive evaluations on a wide range of indoor RGB sequences demonstrate the scalability and predictive power of our approach. Overall, we make the following contributions:
- A novel NeRF-based dense SLAM system that can handle RGB input conditions and allow high-quality novel view synthesis by using depth estimation.
- A ray reparameterization and a Gaussian encoding input that represents the region over which the radiance field should be integrated for neural fields to efficiently represent unbounded scenes.
- We demonstrate state-of-the-art performance on scene reconstruction based on multiple datasets and show competitive results with recent SLAM systems.
2 Related work
Monocular Visual SLAM In the realm of robotic navigation and augmented reality, Simultaneous Localization and Mapping (SLAM) plays a pivotal role. Monocular Visual SLAM has attracted significant attention due to its cost-effectiveness compared to stereo SLAM. This technique constructs the 3D map of the surroundings while simultaneously estimating the camera’s position and orientation within that map. By processing sequential frames of video, monocular SLAM algorithms detect and track visual features across images, using them to calculate depth and structure from motion. Some works focused on using more efficient data structures like VoxelHashing [5, 27, 20] and Octrees [36, 40] to improve performance. The most notable method in this domain is ORB-SLAM3 [4], a versatile and accurate system that utilizes ORB features to achieve robust tracking and mapping under a wide range of conditions. Another innovative approach is DROID-SLAM [34], which introduces deep learning to enhance tracking accuracy and the overall robustness of the SLAM system. While these methods have set high benchmarks, they are primarily reliant on feature extraction and matching, which can be challenging in feature-poor or dynamic environments.
NeRF-based SLAM Recent advances in the Neural Radiance Field have shown powerful ability in scene reconstruction. In this evolutionary sequence, NICE-SLAM [46] and iMAP [30] were the first systems to incorporate these concepts, effectively employing NeRF [15] to render intricate environmental geometry by optimizing neural fields. While they established a strong foundation, the lack of robust loop closure detection became apparent. This results in a susceptibility to drift over time, especially in longer sequences or larger environments where re-localization is critical. GO-SLAM [43] introduces loop closure detection based on the previous work. It determines loop closure by detecting co-visibility between the current and historical keyframes and also attempts to map by relying only on RGB inputs. It has made significant progress in solving the scale drift problem, but the lack of the original depth information of the scene in its RGB input mode leads to a large error in scene scale recovery. Subsequently, HI-SLAM [42] emerged, striving to overcome the flaws inherent in GO-SLAM by implementing a more efficient loop closure and a priori depth estimates from Hessian matrices. However, the systems mentioned above are incapable of mapping in unbounded scenes.
Figure 2: Overview of MoD-SLAM. We demonstrate the monocular mode of our system, which uses RGB streams as its input. 1) Within the tracking module, the system ascertains the camera poses and depth values while executing loop closure based on a co-visibility check. Simultaneously, it selects keyframes that are subsequently fed into the mapping module. 2) The mapping module leverages the RGB map derived from these keyframes to forecast depth values via a depth estimation module and depth distillation process. Concurrently, Gaussian encoding is applied to the keyframes, and the resultant spatial mean and covariance are fed into MLPs.
3 Method
For an input RGB stream, we employ a two-step process: first, we estimate the depth values of keyframes using a depth estimator, and subsequently, we refine these depth estimates using depth distillation to filter inaccuracies (Sec. 3.1). We employ a Gaussian spatial information collection approach within the region to capture both the geometric structure and the visual appearance of the scene. In the case of a scene without distinct boundaries, we employ a contraction function to deal with long-range spatial information (Sec. 3.2). Specifically, we achieve loop closure detection to eliminate scene drift (Sec. 3.3). Finally, employing a NeRF-based volume rendering method (Sec. 3.4), we generate color and depth values for each pixel, facilitating system network optimization (Sec. 3.5). Fig. 2 shows an overview of our method.
3.1 Prior Depth Estimation
Our proposed monocular SLAM system contains a depth estimation module and a depth distillation module. More precisely, the input RGB stream serves dual purposes: first, as a dataset for normal tracking and mapping process, and second, as input to a pre-trained depth estimation module. This module generates depth information for the scene, which is then distilled through a depth distillation module to refine the depth data. The use of priori depth estimation significantly enhances the accuracy of scale in scene reconstructed through monocular mode. Our depth estimation unit is inspired by ZoeDepth [3].
Depth Estimation Fig. 2 shows the components of our depth estimation module. Our depth estimation module contains two parts: relative depth estimation and metric depth estimation. In the relative depth estimation module, we adopt the training criterion of MiDaS [24] and use the DPT decoder-encoder architecture [23] as backbones for feature extraction to predict the relative depths. In the metric depth estimation module, we follow ZoeDepth [3] by proposing the metric bins module to output the metric depth.
The findings from [3, 13] indicate that utilizing a singular metric head across various datasets fails to yield effective predictions. Moreover, we find that predicting depth using only the pre-trained ZoeDepth model leads to scale non-conformity. There is also not a linear multiplicative relationship between the scale predicted by ZoeDepth and the scale predicted by the mapping process. Therefore, we set multiple metric heads in our model and fine-tune ZoeDepth using specific datasets. These strategies constrain the scale of the depth information predicted by our depth estimation module.
Depth Distillation Due to constraints within the dataset, coarse depth annotations, etc., the depth information derived from single-view depth estimation may deviate from actual depth data. Relying directly on the depth information from the estimation module for supervision could lead to inaccuracies in the reconstructed scene. To address this issue, we introduce a depth distillation module. Building on the foundation laid by prior research [37, 18], our module focuses on two key aspects of depth map: spatial continuity and spatial correspondence.
Given a pixel set $\mathbf{P}$ of an RGB image $\mathbf{I}$, we calculate the depth $\mathbf{\widehat{D}}_{k}$ of each pixel point $k_{n}\in\mathbf{P}$. Simultaneously, we extract the depth $\mathbf{D}_{k}$ which shares the same spatial position as $\mathbf{\widehat{D}}_{k}$ from the depth estimation module. We ensure the spatial correspondence of the depth map by introducing a spatial correspondence regularization term. We randomly sample two pixels $k_{m},k_{n}\in\mathbf{P}$, the depth spatial correspondence loss is introduced as:
$\centering\mathcal{L}_{cor}=\frac{1}{M}\sum_{k_{m},k_{n}\in\mathbf{P}}\sum_{\mathbf{D}_{k_{m}}\leq\mathbf{D}_{k_{n}}}\mathbf{max}(\widehat{\mathbf{D}}_{k_{m}}-\widehat{\mathbf{D}}_{k_{n}}+\tau,0),\@add@centering$ (1)
where $M$ represents the count of $\mathbf{D}_{k_{m}}\leq\mathbf{D}_{k_{n}}$ is observed among the total samples taken. $\tau$ is a minimal constant introduced to accommodate slight variations in the depth map. We punish the network if the depth information derived from the depth estimation module lacks spatial correspondence with the depth information obtained through NeRF.
Furthermore, we propose the spatial continuity regularization term to constrain spatial continuity between two depth maps. To achieve it, we randomly sample pixel $k_{m}\in\mathbf{P}$. Then, within $k_{m}$ neighboring 3x3 pixel grids, we extract the depth value of each grid denoted as the set $\mathbf{\Theta}_{k_{m}}$. The depth spatial continuity loss is given by:
$\centering\mathcal{L}_{con}=\frac{1}{M}\sum_{k_{m}:T(k_{m})\geq 4}\sum_{\mathbf{D}_{k_{n}}\in\mathbf{\Theta_{k_{m}}}}\mathbf{max}(|\widehat{\mathbf{D}}_{k_{m}}-\widehat{\mathbf{D}}_{k_{n}}|+\tau^{\prime},0),\@add@centering$ (2)
where $T(k_{m})$ represents the count of $\mathbf{D}_{k_{n}}\in\mathbf{\Theta}_{k_{m}}$, holding $|\mathbf{D}_{k_{m}}-\mathbf{D}_{k_{n}}|<\tau^{\prime}$ true. $\tau^{\prime}$ signifies a minor margin between the current pixel point and its adjacent pixel points. $M$ denotes the aggregate count of samples satisfying the criteria within set $\mathbf{P}$.
3.2 Scene Representation
Multivariate Gaussian Encoding  NeRF constructs the training dataset by projecting a fine ray to the center of each pixel point and then performing coarse and fine sampling. This approach leads to missing information. Inspired by the Mip-NeRF [1], we direct a cone-shaped projection to the pixel rather than slender rays. Projecting the cone allows us to collect more detailed information in the three-dimensional space and transfer it to MLPs for training. The training speed of the model is also greatly improved by eliminating the coarse and fine sampling processes.
To compute the above features, we approximate the conic truncation with a multivariate Gaussian to compute the features in a particular space. The proximity of neighboring sampling points, denoted as $t_{0}$ and $t_{1}$, often results in a minimal distance $t_{1}-t_{0}$, which can compromise stability during division. Consequently, we introduce $t_{\mu}=(t_{0}+t_{1})/2$ and $t_{\delta}=(t_{0}-t_{1})/2$ as refined parameters to bolster the system’s stability. The multivariate Gaussian features are represented by three variables: the mean distance along the ray $\mu_{t}$, the variance along the ray $\sigma_{t}^{2}$, and the variance perpendicular to the ray $\sigma_{r}^{2}$:
$\centering\mu_{t}=t_{\mu}+\frac{2t_{\mu}t_{\delta}^{2}}{3t_{\mu}^{2}+t_{\delta}^{2}},\quad\sigma_{t}^{2}=\frac{t_{\delta}^{2}}{3}-\frac{4t_{\delta}^{4}(12t_{\mu}^{2}-t_{\delta}^{2})}{15(3t_{\mu}^{2}+t_{\delta}^{2})^{2}},\@add@centering$ (3)
$\centering\sigma_{r}^{2}=r^{2}(\frac{t_{\delta}^{2}}{4}+\frac{5t_{\delta}^{2}}{12}-\frac{4t_{\delta}^{4}}{15(3t_{\mu}^{2}+t_{\delta}^{2})}),\@add@centering$ (4)
r is the perpendicular distance between the sampling plane where the ray is located and the ray at the center of the pixel point after transformation by spherical coordinates. Define the position of the origin of the camera in the world coordinate system as $\mu=\mathbf{o}+\mu_{t}\mathbf{d}$. All of the above are functional relationships obtained in the camera coordinate system. To determine the position of the pixel point in the world coordinate system more rationally, we transform the mean and covariance of the variables expressing the multivariate Gaussian from the camera coordinate system to the world coordinate system as follows:
$\centering\sigma=\mathbf{d}\mu_{t},\quad\mathbf{\Sigma}=\mu_{r}(\mathbf{I}-\mathbf{d}\mathbf{d^{T}})+\mu_{t}\mathbf{d}\mathbf{d^{T}},\@add@centering$ (5)
following the above equations, we can obtain a multivariate Gaussian representation for the position as well as the direction in a particular space.
Figure 3: Scene Reparameterization. To manage unbounded scenes, we use a contract function to map mean and covariance data from spatial sampling (black dashed line) directly to a new range (red solid line). We save the data within the sphere of radius 1 (yellow region) while mapping the data from regions beyond a radius of 1 into the spherical coordinate system with a radius of 2 (orange region).
Unbounded Scenes Previous research [2] has demonstrated that the predictive efficacy of MLPs significantly diminishes when applied to datasets characterized by large disparities. Thus, for bounded scenes, a normalization method is generally used to restrict all the data in the training set to a specific region to enhance the predictive capability of MLPs. Nevertheless, for scenes without defined boundaries, normalization of the scene is impracticable owing to the lack of essential boundary scale information. Therefore, we define $f(\mathbf{x})$ to reparameterize the data as follows:
$\centering f(\mathbf{x})=\lambda\mathbf{x}+(1-\lambda)(2-\frac{1}{\lVert\mathbf{x}\rVert})(\frac{\mathbf{x}}{\lVert\mathbf{x}\rVert}),\quad\lambda=\begin{cases}1\quad\lVert\mathbf{x}\rVert\leq 1\\ 0\quad\lVert\mathbf{x}\rVert>1\end{cases}\@add@centering$ (6)
This can be interpreted as an extension of the normalized device coordinate (NDC). We make the scene shrink along the inverse depth direction, thus enabling the network to train on and interpret unbounded geometric scenes. However, since in Equ. 5 above we used the mean and covariance to represent information about the cross-section within the conic, we also need to map the mean and covariance to it.
To achieve this, first, we linearize $f(\mathbf{x})$ by performing a first-order Taylor expansion at the point $\mathbf{\mu}$:
$\centering f(\mathbf{x})=f(\mathbf{\mu})+\mathbf{J}_{f}(\mathbf{\mu})(\mathbf{x}-\mathbf{\mu}),\@add@centering$ (7)
where $\mathbf{J}_{f}(\mathbf{\mu})$ is the jacobian of $f$ at $\mathbf{\mu}$. After doing that, we can obtain mean function $E(\mathbf{x})$ and covariance function $D(\mathbf{x})$. our choice for $f$ is relied on Equ. 6.
$\centering f(E(\mathbf{x}),D(\mathbf{x}))=\begin{cases}E(\mathbf{x}):f(\mathbf{\mu})\\ D(\mathbf{x}):\mathbf{J}_{f}(\mathbf{\mu})D(\mathbf{x})\mathbf{J}_{f}(\mathbf{\mu})^{T}.\end{cases}\@add@centering$ (8)
3.3 Loop Closure
SLAM calculates the camera’s position for the next frame based on the position of the previous frame, resulting in the accumulation of errors. Loop closure detection is introduced to eliminate the drift. Our system is divided into four stages: loop closure candidate frame detection, $sim3$ change calculation, loop closure fusion and pose optimization. For the current key frame $K_{i}$, we find the best three loop closure candidates and fusion candidates $K_{m}$ based on the similarity. Then for each candidate frame, we define a localized window containing 5 keyframes with the highest covariance with the candidate frame and the map points. To avoid false positives, the loop closure candidate frames are further subjected to covariance geometry checking as well as timing geometry checking. After that, the loop closure candidate frames selected after verification are used to construct a local window with the current keyframes for map fusion. Finally, the keyframes are optimized to eliminate the drift during tracking.
3.4 Depth and Color Rendering
Inspired by the recent work in NeRF, we propose to also use Neural radiance field to reconstruct 3D scenes.
Given the current camera pose and camera intrinsic matrix, we can calculate the viewing direction r and change it to the world coordinate. Unlike NeRF, since our system is also used to process borderless scenes at the same time, we propose here a more efficient sampling technique, the differential sampling technique. This sampling technique aims to sample more in the near part of the ray and sparsely in the far part. To accomplish this idea, we first sample the ray at the proximal $n$ as well as at the distal $f$, and then obtain a series of sampling points by sampling uniformly in this space $\{s_{1},...,s_{N}\}$. After that, we map the sampling points to the following function:
$\centering s^{\prime}=\frac{1}{s\frac{1}{f}+(1-s)\frac{1}{n}},\@add@centering$ (9)
and then we get the new sampling points as $\{s_{1}^{\prime},...,s_{N}^{\prime}\}$.
To complete the scene reconstruction, we randomly extract $N$ rays and $M$ pixel points in the current frame and keyframes. We input the position set $\mathbf{P}$ to the depth prediction module in Sec. 3.1 to get the true depth $\mathbf{D}$. We do a multivariate Gaussian variation of the pose set $\mathbf{G}$ in Sec. 3.2 to get the mean as well as the covariance set. The mean and covariance are then passing into the MLPs to obtain the predicted volume density as well as the predicted color. Volume density is used to calculate the occupancy probability $o_{i}$ and the ray termination probability at each sample can then be calculated as $w_{i}=o_{i}\prod_{i=j}^{i-1}(1-o_{j})$. Eventually, the depth and color can be rendered as:
$\centering\widehat{\mathbf{D}}=\sum_{k=1}^{N}w_{i}d_{i},\quad\widehat{\mathbf{C}}=\sum_{k=1}^{N}w_{i}c_{i},\@add@centering$ (10)
where $d_{i}$ is the corresponding depth values of sampling points and $c_{i}$ is the rgb values of sampling points.
3.5 Tracking and Mapping
Camera Tracking For more accurate position tracking, in this work, we employ DROID-SLAM [34] as the front end of our system due to its superior position prediction capabilities and minimal position drift. We take the RGB stream as input, and for each frame maintain two state variables: camera position and inverse depth. The pose set $\left\{\mathbf{G}_{t}\right\}^{N}_{t=0}$ and the inverse depth set $\left\{\mathbf{d}_{t}\right\}^{N}_{t=0}$ are unknown state variables, which are iteratively updated during the inference process of processing a new frame. Each input image is processed through a feature extraction network with a context network.
In the network training phase, we train our network by using a combination of pose loss and depth loss. We pre-compute a distance matrix of size $\mathbf{K}*\mathbf{K}$, where $\mathbf{K}$ is the length of each input image. This matrix is used to store the size of the average optical flow between each pair of frames. Then, we compute the depth loss $\mathcal{L}_{d}$ generated by the predicted depth and the true depth, and the position pose loss $\mathcal{L}_{v}$ generated by the predicted bit pose and the true position pose. The loss is taken as the average l2 distance between the two optical flow fields [41]:
$\centering\mathcal{L}_{tracking}=\mathcal{L}_{v}+\lambda_{v}\mathcal{L}_{d},\@add@centering$ (11)
where $\lambda_{v}$ is a weight used to limit the depth loss to the overall tracking loss. In this system, we take $\lambda_{v}$ equals to 0.1.
Mapping To enhance the fidelity of the reconstructed scene to the actual scene, we initially sample $M$ pixels from the keyframes. Subsequently, the network can be optimized by minimizing the loss.
We use the method mentioned in [16] to render the color of the pixel point $\widehat{\mathbf{C}}$. To optimize the network, we minimize the loss between the predicted color and the true color:
$\centering\mathcal{L}_{rgb}=\frac{1}{M}\sum_{m=1}^{M}|\widehat{\mathbf{C}}_{m}-\mathbf{C}_{m}|,\@add@centering$ (12)
We also use Equ. 1 and Equ. 2 to constrain depth value to supervise the reconstruction of this system and better recover the scale:
$\centering\mathcal{L}_{depth}=\mathcal{L}_{cor}+\lambda_{con}\mathcal{L}_{con},\@add@centering$ (13)
where $\lambda_{con}$ denotes the hyperparameter employed to modulate the weight of $\mathcal{L}_{depth}$. Furthermore, following [2], we also introduce distortion loss, which is used to eliminate artifacts in pictures caused by discontinuities in volume density over tiny areas. Distortion loss is defined as:
$\centering\mathcal{L}_{dist}=\sum_{i,j}w_{i}w_{j}\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\rvert+\sum_{i}\frac{1}{3}w_{i}^{2}(s_{i+1}-s{i}),\@add@centering$ (14)
where the first term minimizes the weighted distance between pairs of points in all intervals and the second term minimizes the weighted size of each individual interval. Additionally, following [10], regularization to the predicted SDF is also introduced in this system. At each iteration, we can get the predicted signed distance of given batch points as: $\widehat{\xi}=f(s;\theta)$. To encourage valid SDF values, especially in unsupervised areas, we adopt the Eikonal term:
$\centering\mathcal{L}_{eik}=\frac{1}{MN_{ray}}\sum_{m=1}^{M}\sum_{i\in N}(1-\lVert\bigtriangledown\mathbf{\phi(s_{i})}\rVert)^{2}.\@add@centering$ (15)
For the SDF, we approximate the SDF values by calculating the depth values along the observation direction as: $\mathbf{b(s)}=\widehat{\mathbf{D}}-\mathbf{D}$. We then define the adopted loss function by its distance from the surface. For $|\mathbf{\phi(s)}|\leq|\mathbf{b(s)}|$,$\forall\mathbf{s}$, the near-surface points($\widehat{\mathbf{D}}-\mathbf{D}\leq\zeta$), following [21, 39] the SDF loss is:
$\centering\mathcal{L}_{sdf}=\frac{1}{MN_{ray}}\sum_{m=1}^{M}\sum_{i\in N}\lvert\mathbf{\phi(s_{i})}-\mathbf{b(s_{i})}\rvert,\@add@centering$ (16)
for the other situation, the far-surface points($\widehat{\mathbf{D}}-\mathbf{D}>\zeta$), the SDF loss is:
$\centering\mathcal{L}_{sdf}=\frac{1}{MN_{ray}}\sum_{m=1}^{M}\sum_{i\in N}max(0,e^{-\alpha\mathbf{\phi}({\mathbf{s_{i}})}},\mathbf{\phi}({\mathbf{s_{i}})}-\mathbf{b}({\mathbf{s_{i}})}),\@add@centering$ (17)
where we set $\alpha$ equal to 5 in this system. In our approach, the treatment of SDF predictions varies based on their values: no penalty is imposed when the prediction is positive and falls below a specified bound; a linear loss is introduced for larger SDF predictions; and, in cases where the prediction is negative, an exponential penalty is applied to ensure smooth transitions between these states.
Our real-time mapping thread optimizes the predicted neural networks by continuously feeding keyframes to achieve ideal 3D reconstruction results. The total loss of the system is:
$\centering\mathcal{L}_{total}=\lambda_{c}\mathcal{L}_{c}+\lambda_{dep}\mathcal{L}_{dep}+\lambda_{dist}\mathcal{L}_{dist}+\lambda_{eik}\mathcal{L}_{eik}+\lambda_{sdf}\mathcal{L}_{sdf},\@add@centering$ (18)
where $\lambda_{c}$, $\lambda_{depth}$, $\lambda_{dist}$, $\lambda_{eik}$, $\lambda_{sdf}$ are the hyperparameters for each loss, controlling the weight of each loss in the total loss.
Figure 4: Reconstruction results on Replica [28]. Compared with baselines, the 3D scenes reconstructed by our method achieve better geometric structure and texture features.
Figure 5: Reconstruction results on ScanNet [7]. We show the full view of the reconstructed 3D scenes. Our system can better reconstruct the scale of the scenes.
4 Experiments
We evaluate our models in both synthetic and real-world datasets, including Replica [28] and ScanNet [7]. We also conducted a comparative analysis of the reconstruction quality of our proposed system against state-of-the-art SLAM systems.
4.1 Experimental Setup
Figure 6: Reconstruction results on ScanNet [7]. Compared with baselines, our system demonstrates enhanced capability in scale recovery and void interpolation within the scene.
Datasets For our quantitative analysis, we implement our system on the Replica dataset and ScanNet dataset. We follow the same pre-processing step for datasets as in NICE-SLAM [46]. Each dataset presents challenging camera dynamics, effectively showcasing our model’s performance. We concurrently utilize these datasets to assess our model in both monocular RGB and RGB-D modes.
Metrics In evaluating camera tracking accuracy, we follow standard SLAM metrics. This involves comparing the trajectory of keyframes against the true trajectory and computing the Root Mean Square Error (RMSE) of the Absolute Trajectory Error (ATE) [29]. Regarding reconstruction quality, we evaluate it by accuracy [cm], PSNR [9], SSIM and F-score [$<$ 5cm %] [30]. In addition, we evaluate the L1 loss [46] between the reconstructed and real meshes. All our experimental results are averaged after three runs.
Implementation Details We run our system on a PC with a 4.70 GHz AMD 7900X CPU and NVIDIA RTX 4090 GPU. We use the pre-trained model from DROID-SLAM [34] for tracking and use the same parameters in the tracking process as it. In the reconstruction process, we choose to sample a ray at several points of $N_{strat}=20$ and $N_{imp}=40$. For loss weights, we choose $\lambda_{c}=1.0$, $\lambda_{dep}=1.0$, $\lambda_{dist}=0.002$, $\lambda_{eik}=0.15$, $\lambda_{sdf}=1.0$. The compared baseline includes classical SLAM systems ORB-SLAM3 [4] and DROID-SLAM [30], newest SLAM systems NICE-SLAM [46], HI-SLAM [42] and GO-SLAM [43] and 3D Gaussian-based SLAM Photo-SLAM [11].
[…]
NICE-SLAM[46] $1.95$ $3.53$ $2.85$ $26.31$ 0.847 $0.3$
RGB-D Go-SLAM[43] 0.34 3.38 2.50 27.38 0.851 8
Ours 0.33 3.11 2.13 30.95 0.882 6
DROID-SLAM[34] 0.42 17.59 5.03 7.68 0.191 21
NICER-SLAM[45] 1.88 x 3.65 25.41 0.827 6
Mono GO-SLAM[43] 0.39 4.39 3.81 22.13 0.733 8
Photo-SLAM[11] 1.09 x x 33.30 0.930 42
Ours 0.35 3.23 2.48 28.37 0.850 8
Table 1: Reconstruction quality and run-time comparison on Replica [28]. Our results were obtained by averaging three runs across seven scenes from the Replica dataset. ’-’ indicates the inability to acquire pertinent data. ’x’ denotes data that has yet to be released.
Scene ID $0$ 0059 0106 0169 0181 0207 0233 Avg.
DROID-SLAM[46] $5.48$ $9.00$ $7.86$ $7.41$ 5.43 $72.23$ $6.76$ $16.31$
Mono Go-SLAM[43] 5.94 8.27 8.42 8.29 7.48 5.31 8.07 7.40
Ours 5.39 7.78 7.64 6.79 4.37 5.63 6.58 6.32
NICE-SLAM[46] 8.64 12.25 10.28 12.93 5.59 10.53 8.09 9.76
GO-SLAM[43] 5.35 7.52 7.74 6.84 3.47 4.78 7.03 6.10
RGB-D Co-SLAM[38] 7.13 11.14 5.90 11.81 7.14 8.75 9.36 8.75
E-SLAM[12] 7.32 8.57 6.53 9.02 5.74 7.96 7.51 7.52
Ours 5.27 7.44 6.73 6.48 3.29 5.31 6.14 5.81
Table 2: ATE RMSE [cm] on the ScanNet dataset [7]. The performance metrics reported by our system for each scene are derived from the average of three executions per scene.
4.2 Results and Evaluation
Evaluation on Replica [28] Table 1 shows the results of the reconstruction of systems on the Replica dataset. Compared to the systems mentioned in the table, our system exhibits superior tracking capabilities in both monocular and RGB-D modes. In comparison with NeRF-based SLAM [46, 43, 45], our approach achieves enhanced reconstruction fidelity. Compared to 3D Gaussian-based SLAM systems [11], our system falls in terms of scene reconstruction accuracy. However, our system outperforms in tracking performance and is more effective in recovering the scale of the scene. Fig. 4 presents a comparative analysis of scene reconstructions by our system against NICE-SLAM [46] and GO-SLAM [43]. Our method demonstrates superior performance that shows smoother scene reconstruction.
Evaluation on ScanNet [7] We evaluated our system across seven distinct scenes in ScanNet dataset [7]. Table 2 presents the tracking results. Our system outperforms others in tracking accuracy, particularly in the large indoor scenes of ScanNet. As illustrated in Fig. 6 and Fig. 5, our system outmatches GO-SLAM [43] by rendering the scene with enhanced detail fidelity, more accurately approximating the real-world geometry and ensuring smoother scene transitions. Furthermore, our framework possesses a pronounced proficiency in addressing scene sparsity through effective void interpolation.
5 Conclusion
We presented MoD-SLAM, a monocular implicit neural-based dense mapping method that demonstrates state-of-the-art accuracy in monocular reconstruction. Our experiments show that by integrating a monocular depth estimation module, our system not only achieves a more accurate scene-scale reconstruction but also excels in reconstructing the details and textures of geometric objects compared to previous SLAM systems. Additionally, we show the superior performance of our system in reconstructing unbounded scenes after employing Gaussian encoding and reparametrization techniques.
Limitations The depth estimation module relies on a pre-trained model; thus, for complex scenes or those significantly divergent from our provided pre-trained scenes, it’s necessary to retrain the depth estimation module. Moreover, compared to traditional SLAM, our method lacks an IMU [32] module, leading to tracking errors. Finally, the integration of depth estimation in mapping under monocular mode results in a slower mapping speed.
Appendix 0.A Appendix
0.A.1 Ablation Study
In this section, we evaluate the choice of our depth estimation module, multivariate Gaussian encoding and reparameterization.
Depth Estimation We investigate the effect of the depth prediction module and depth distillation on the performance of our system. Table 5 shows the reconstruction performance of the system in monocular mode with the depth estimation module removed and the depth estimation module incorporated. Without the depth estimation module, our system suffers a dramatic drop in both tracking and mapping performance. However, with the cooperation of the depth estimation module and depth distillation module, the monocular mode performance of our system closely approaches that of the RGB-D mode.
Losses We further evaluate the impact of each loss on the reconstruction performance of the system. Table 3 shows the impact of each loss on the reconstruction performance of the system, and it’s easy to find that only by relying on $\mathcal{L}_{c}$ and $\mathcal{L}_{depth}$ cannot lead to the best reconstruction result. $\mathcal{L}_{sdf}$ has a more significant impact on the improvement of reconstruction quality because it affects the expression and reconstruction of the object’s geometric surface information. In addition, $\mathcal{L}_{eik}$ and $\mathcal{L}_{dist}$ also have a positive impact on improving the reconstruction performance, one of them is used to optimize the geometric surface and the other is used to eliminate artifacts. Integrating all terms produces the best result.
Unbounded Scene We also conduct tests to assess the impact of reparameterization Equ. 6 and Gaussian encoding Equ. 5 on reconstructing scenes with no defined boundaries. For comparison, we used NEWTON [14] as a baseline to evaluate our system’s performance. The quantitative data in Table 5 indicates that our reparameterization and Gaussian encoding significantly improve the accuracy in both scaling and geometric reconstruction of unbounded scenes. These approaches notably enhance our system’s overall performance. Our data were acquired using the RGB-D mode of our system within the ScanNet 0207 scene. NEWTON’s data are from the original paper [14]. The results on the ScanNet [7] indicate that our model shows greater robustness and more precise mapping abilities compared to NEWTON.
Figure 7: (a) shows the reconstruction result on Replica [28]. (b) shows the run time and GPU memory used in SLAM systems.
[…]
1 0 0 0 0 34.33 15.78
1 1 0 0 0 85.79 27.76
1 1 1 0 0 85.98 27.89
1 1 1 1 0 89.01 28.13
1 1 1 1 1 89.73 28.37
Table 3: The impact of various losses. All data are averaged over three runs in monocular mode on our system under seven scenes of the Replica dataset. ’1’ represents use this loss while ’0’ means ignore.
Table 4: Depth ablation study. We explore the effects of the depth estimation module (DepE.) and the depth distillation module (DepD.) in our system.
ATE[cm]$\downarrow$ F-$\uparrow$
None 0.46 83.79
DepE. 0.39 87.96
DepE./DepD. 0.35 89.43
RGB-D 0.33 90.96
[…]
Con. 23.85 0.749
Con-Gau. 25.02 0.787
None - -
NEWTON[14] 23.15 0.702
Table 4: Depth ablation study. We explore the effects of the depth estimation module (DepE.) and the depth distillation module (DepD.) in our system.
Table 5: Unbounded scene ablation study. We investigate the effects of the constructed function (Con.) and Gaussian encoding (Gau.) in our system.
0.A.2 Runtime Analysis
We test the tracking and the mapping time of NICER-SLAM [45], GO-SLAM [43] and our system using the same number of sampling points. Fig. 7(b) shows that our system is faster than GO-SLAM and NICER-SLAM in the tracking process. However, because of the depth estimation module, our system is slightly slower than GO-SLAM in the mapping process, but higher than NICER-SLAM. In addition, Fig. 7(b) also demonstrates that our model size is smaller and uses less memory. GO-SLAM and our system are benchmarked with an NVIDIA RTX 4090 GPU while NICER-SLAM is benchmarked with an A100.
Scene fr1-desk fr2-xyz fr3-office
GO-SLAM[43] 1.500 0.600 1.000
RGB-D Co-SLAM[38] 2.400 1.700 2.400
Ours 1.382 0.641 0.939
DROID-SLAM[34] 2.713 1.304 2.573
DeepV2D[33] 21.391 6.950 11.922
Mono ORB-SLAM3[4] 1.534 0.720 1.400
GO-SLAM[43] 2.381 1.093 1.773
Ours 1.481 0.686 1.101
Table 6: ATE RMSE [cm] on the TUM dataset [29]. The performance metrics reported by our system for each scene are derived from the average of three executions per scene.
0.A.3 Additional Experimental Results
TUM Dataset [29] We further test the tracking accuracy on the TUM dataset [29]. As shown in Table 6, Our method not only achieve the most accurate tracking performance in RGBD mode, but also outperforms classic SLAM systems as well as GO-SLAM [43] in monocular mode.
Replica Dataset [29] Here we show more results on Replica [29]. Fig. 8, Fig. 9 and Fig. 7(a) demonstrate the superior scene reconstruction capability of our proposed method in monocular and RGB-D mode, and also exhibiting enhanced smoothness in comparison to GO-SLAM [43].
Figure 8: Reconstruction results on Replica [28]. Qualitative comparison on Replica office4 and room2.
Figure 9: Reconstruction results on Replica [28]. Qualitative comparison on Replica office1 and office3.